[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Jun 23, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Thu Jun 23 13:07:57 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
Lightning data is not coming in to N-AWIPS. Matt will investigate.
HSU:
TSB is still working on restoring processing that was lost due to the disk  
crash on Mike's workstation. Matt Sardi successfully migrated the ntrans  
processing to one of the operational compute servers, including the  
soundings. The remaining applications to be recovered are the sounding time  
section processing for the NHC web page, and the creation of the model  
products that are viewed via the seahorse server. The disk drive was sent  
out to a disk recovery service to speed up the restoration of those  
applications.

Matt worked with NCF to find the source of the spurious pre-TCM and TCM  
messages. These were not actually sent out from NHC, but where printed when  
they were opened on a local NHC AWIPS system.
Other items:
Salim Leyva will be at NHC starting today along with an NCO contractor to  
help restore the local GOES-east ingest system. They will also assist with  
installing the new CARCAH UPS system. Salim will train the NCO contractor  
to assist with the satellite ingest in the future.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160623/2c5bc78c/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list